OpenStudy (anonymous):

For 3 Weeks , Jessica worked 2 hours per day, 4 days per week baby sitting. If she earned 3.50 per hour, how much did she earn in all ? Write your answer with a doller sign and include 2 decimal places since its money.????

OpenStudy (anonymous):

Also one more Maria and paul collect marbles. Paul has one third as many marbles as maria. Together they have 48 marbles . How many marbles does Maria have??

OpenStudy (anonymous):

Total hrs worked by Jessica=3*4*2=24 hrs for 1 hrs she getting 3.50 $ hence total earning by Jessica=3.5*24=84.00$

OpenStudy (anonymous):

Also one more Maria and paul collect marbles. Paul has one third as many marbles as maria. Together they have 48 marbles . How many marbles does Maria have?